Blue side video board noise (2009-March-30)
- The video processor board on the left most CCD of LRIS Blue
failed and was replaced with a spare. The spare board
exhibits low level noise as seen in this
The stddev in the bias is 2.7 in the noisy amp relative to
2.1 for the right most amplifiers. We are in the process of
repairing or replacing the old video board. The failed board
was sent to the manufacturer for assessment.
Red Side Halos Problem fixed (2009-March-20)
- We removed the LRIS red camera and found some condensation
on both sides of the dewar window. The dewar window was
cleaned, and the dewar was pumped out. Following cool down,
the halos reported on 2009-Feb-15 were no-longer detected.
Starlists Now Allow PA Info (2009-Feb-20)
- For years, observers have been forced to provide their
position angle information to the OA verbally. The new starlist
file format now permits you to enter the position angle
for your target using the ROTMODE and
ROTDEST keywords. Here is an example of how to
specify the position angle:
# slitmask with specified rotator position
PA_targ_name 00 24 00.00 +16 00 00.00 2000.0 rotdest=123.45 rotmode=pa
# target observed at parallactic angle
VA_targ_name 12 34 56.78 +01 23 45.67 2000.0 rotdest=0.0 rotmode=vertical
Red Side Halos (2009-Feb-15) -Fixed 2009 March 20
- Images of point sources on the red side only have halos at
a level of 1% of the source peak flux. An
image of the grid
of holes mask shows clearly this effect. This effect is also
present in spectra on the red side. We suspect at this time
that condensation is present on the red side dewar
window. The Red side camera will be removed in early March
to address this problem.
New Slit Guider Camera commissioned. (2009-Feb-15)
- A new slit guiding camera was commissioned for use during
engineering nights on 12 and 14 Feb. The new slit guider is
rotated such that the slit is 29.5 degrees relative to the
columns on the detector. The new camera is sensitive enough
to reach a signal-to-noise of 10 in 1s on a star of
v-mag=19.5. The unvignetted FOV is expected to be 1.6 arcmin
diameter, and at 40% vignetting, the FOV is expected to be
2.2 arcmin. For observers who use pickoff mirrors, the new
default pixel location on the guide camera for a pickoff
mirror guide star is 459 618. A new slitb pointing origin
was also defined to move the star away from a bad column on
the red side detector. An image of M13 is available
here. The dark line
going through the field is a 1.0 arcsec wide long slit.
New Hatch Commissioned but has light leaks (2009-Feb-15)
- All observers are advised to acquire
internal calibrations after the dome lights are turned off
because the hatch has light leaks.
As part of the new guider upgrade, a new hatch was
needed to house all the new guider optics.
The new hatch was tested during engineering nights
on 12 and 14 Feb, and it was determined that it has light
leaks. An internal G filter pseudo-flat acquired with all
internal lamps off, hatch closed, and dome lights on
revealed the light leaks seen in the image
here. The peak
flux from the light leaks is ~3000 ADU, and it varies as a
function of rotator position, telescope elevation, and
telescope azimuth. This problem will be address in either
early or late March.
831/8200 Grating available (2009-Jan-22)
- We received and commissioned a new 831/8200 grating. The
grating is available for general use, and there is an option to
select the grating in the LRIS configuration form.
LRIS slit guider upgraded (2008-Dec-23)
- We have commissioned a new slit guide camera and
electronics for LRIS. This was an emergency commissioning of a
new guider camera because the old slit guider camera was flooded
with glycol after a coolant connector seal failed. The old
camera was a Photometrics CH250. The new camera is a
Photometrics CH350. New electronics are part of the
commissioning and the electronics are not inside the electronics
bay. The new camera has a wider FOV with significant vignetting
in the corners (see image). It
is also a more sensitive camera with a higher gain and less
readnoise than the old camera by a factors of 3 and 4,
respectively. With the old guider, objects were centered at
pixel 191, 143. Last night's commissioning result show that
objects should now be centered at pixel 277, 290. In the MAGIQ
software, Observing Assistants should see two cameras: one for
LRIS lris50m (as before), and another called lrisSlit for the
slit-guider. Both MAGIQ and XGUIDE were updated and released for
831/8200 Grating status (2008-Nov-13)
- A new 831/8200 grating was ordered and is scheduled to
arrive in January 2009. It was determined that the old
grating could not be used inside LRIS due to the risk of
additional fracturing. The new grating will be made from the
same master grating as the original LRIS 831/8200 grating.
Desktop Menu Returns (2008-Nov-13)
- The desktop menu or background menu is now used to run the
instrument, replacing the LRIS workspace menu that appeared
in the upper left corner of every display.
New IRAF tasks make LRIS mask alignment more convenient (2008-Nov-03)
- Three new tasks are now available in the
keck.lris package to assist observers with aligning
All three tasks presume that the file listing the box
coordinates is named slitmask.box, where
slitmask is the name of the slitmask as shown in the
- do_check_boxesb will wait until the image is
written (if one is in progress), then will read the image name and run
check_boxesb on it.
- do_xboxb will wait until the image is
written (if one is in progress), then will read the image
name and run xboxb on it.
- get_seeing will estimate the current seeing
by measuring the profiles of the stars in the boxes on a
slitmask alignment image.
Grating 831/8200 damaged and unavailable (2008-Oct-9)
- The 831/8200 grating is damaged and has been temporarily
taken out of service. Three fragments, each roughly one
cubic inch in size, were found in the corners of the
grating. We are in the process of assessing the grating
substrate to determine the risk of future fragmentation, and
we are exploring options for repairing and replacing the
grating. If we determine that the only solution is to
replace the grating, the time frame for replacing it is 10
weeks. If we determine that the damage will not impact
science and there is no risk of future fragmentation, then
the time frame for repair is 6 weeks.
Web page makeover (2008-Aug-19)
- All LRIS web documentation was updated to the new instrument page
format. If you find content or pages missing, please contact
your support astronomer.
New I band filter installed (2008-Jul-11)
- A new I band filter is available for use with LRIS. This
filter is an interference type filter with a relatively
square passband with FWHM from 6986-8204 Angstroms. Please
filters on-line documentation to view the transmission
NB4000 filter has a long-wavelength leak (2008-Jul-10)
- Observers noted that the NB4000 filter has developed a
long-wavelength leak past roughly 7000 Angstroms. Engineering
data of this filter was acquired on 8 July 2008 to
characterize the long wavelength leak. Spectroscopic flats
of the internal halogen lamp using the 1 arcsecond long-slit
and 150/7500 grating tuned to 7000 angstroms were acquired
with and without the filter in beam. The filter spectrum was
divided by the flat to produce a relative transmission
spectrum. Please see
to view the plot.
New MAGIQ Guider Software released (2008-Jul-7)
We are in the process of upgrading the acquisition and guide
cameras on all WMKO instruments, and LRIS has been selected
as the next instrument to receive a hardware upgrade.
An important element of our upgrade strategy is to first
deploy the MAGIQ software for operation with the two
existing LRIS guide cameras. The new software provides all
of the functionality of the existing acquisition and guiding
software while improving routine tasks such as star list
management, field identification, acquisition, and guiding.
Successful testing of MAGIQ software with the existing slit
viewing and offset guiders was completed on May 27 and June
27, respectively. The MAGIQ software is released for use,
and you are encouraged to use MAGIQ during your observing
The hardware portion of the guider upgrade will take place
at a later date.
Slitmask Design Software (2008-Apr-18)
At this time, it is recommended that observers use the
Autoslit software to design slitmasks for LRIS.
Dr. Arjun Dey and his observing team compared the
mask designs that were output from both autoslit and the
ucsc lris package and found significant difference in the
slit locations. The specific recommendation by the author
of the UCSCLRIS package is that observers use the AUTOSLIT
field mapping until the UCSCLRIS software is updated.
FVWM with LRIS (2007-Aug-29)
FVWM is a multiple virtual desktop window manager for the X
windows system which is now in use with all LRIS
accounts. With the switch from open-windows to FVWM, a couple
of other changes were required. Most instrument operations
are now found on the "LRIS workspace menu" which is
automatically launched on startup and will be located in the
upper left corner of each display. The "LRIS workspace menu"
replaces the open-windows background menu, and is similar to
the "DEIMOS workspace menu." Second,
clicking on the virtual desktop will bring up a FVWM
background menu which is used mostly to start VNC, start
local xterms, start browsers, and instrument menus. The FVWM
is not used to run the LRIS software, and there are no
options provided on that menu to run the instrument. The web
documentation has been updated with these changes, but
please let your SA know of any inconsistencies.
Blue Side compass rose (2007-Aug-29)
A compass rose that works with the blue side figdisp
display is released for observer use.
Grating offsets applied (2007-Aug-08)
Grating offsets are now in use so that the actual
central wavelength is closer to the desired central
wavelength. Each grating and port combination requires a
different offset. All gratings in all ports have been
mapped. For most grating and port combinations, the
accuracy to which the grating is positioned is better than
10 angstroms. Gratings in port 3 may exhibit larger
offsets of up to 50 A. In the past, observers' noted
offsets from 100 to 1000 Angstroms. If you notice a large
offsets of > 20 Angstroms in the desired central wavelength compared to the
actual central wavelength, please inform you support
astronomer and add a comment to your post observing
I filter broke. (2007-July-20)
On 18 July, the I filter was found broken inside the
instrument. One of the largest pieces was lodged in the
grism mechanism and was preventing movement of the grism
mechanism. There is a 3-4 mm gouge in the back surface of
the 600 line grism. Fortunately, the gouge is located at
the very edge of the grism and this should not alter the
grisms optical performance. Because the filter pieces
tumbled inside LRIS for at least half a night, we are
taking the time to inspect all optical surfaces that may
have been exposed to filter fragments.
In the long term, we expect to have LRIS ready for
operation by the start of the next LRIS observing session
which begins 8 Aug. LRIS will be without an I band filter
on the red side until a suitable replacement is acquired.
ADC and LRIS together at last (2007-July-09)
Please see the ADC docs.
LRIS-ADC MODE: The default observing mode for LRIS as of semester
2007B is to use LRIS with the new Keck I Atmospheric Dispersion
Compensator (ADC), regardless of what you requested on your observing
proposal. The presence of the ADC changes the focal plane scale of
LRIS by a small but measureable amount, meaning that:
(a) old slitmasks designed for use without the ADC should not be
used with the ADC, and
(b) new slitmasks must be designed with software that accounts for
If you plan to use LRIS slitmasks, then you must download the new,
ADC-compatible version of Judy Cohen's AUTOSLIT software (v3.09) or
the similar FABMASK package by Drew Phillips. Please contact your
support astronomer for ADC-related questions.
LRIS Blue side shutter fixed, and CCD readout problem
- The blue side shutter was fixed between 31 April and May
5. A bushing on the shutter was worn. and galling in the
bushing was causing the shutter to stick. A dry lube was
applied to the bushing.
The CCD can be read-out using all four amplifiers
following a repair of the CCD signal cable.
The ccd readout problem experienced was that the second
amplifier did not see the signal. This meant that the
detector had to be read-out in single amp mode, which is
slower. During the shutter repair, technicians discovered
that the CCD cable
had a busted wire. Re-soldering the wire to the connector
has fixed the readout problem.
LRIS Blue side CCD readout problem (2007-April-18)
- As of April 17, one of the amplifiers on the "left" CCD of
the blue-side mosaic is failing to read out properly.
Troubleshooting suggests that the problem is a bad cable between
the dewar and the saddlebag. Since we have no spare for the
cable, we are unable to repair this fault until the run is over.
In the meantime, the workaround is to read out both blue CCDs
using the "right" amplifiers only. The command to put the blue
side into this mode is:
Readout times will be double the usual for two-amp mode on the
LRIS Blue side shutter is stuck open (2007-April-16)
- The LRIS blue channel shutter has failed again, and work to
replace the shutter will begin on 23 April 2007. The LRIS blue
channel shutter should be available in time for the next run which
starts 7 May 2007.
For now, the trapdoor may be used as
a shutter for the blue side.In this mode, red side data may still be acquired
simultaneously. Your SA will review the scripts with you
upon your arrival, but please see
trapdoor_as_shutter.html for a brief description of
the observing scripts.
Observers who used LRIS during April 2007 should
carefully examine blue side data to ensure the shutter
was not stuck open during their observations.
LRIS Red channel missing columns (2007-April-13)
LRIS Troughput Measurements (2007-Mar-12)
- To regularly monitor the throughput of LRIS, observers will be asked
for permission to take spectrophotometric observations of a designated
standard star during twilight. A script called observe_flux_
standard: verifies that the telescope is pointed to a known
LRIS flux standard star by checking DCS keyword TARGNAME; checks
the current grating and build appropriate exposure list; saves
current LRIS configuration;
checks that SKYPA is correct for this star; sets the correct
pointing origin (SLITB); configures for slitless spectroscopy;
acquires exposures; and restores the observer's configuration.
Before running the script the observer should be set up to
acquire spectroscopic measurements on the red side (blue side
optional) and the OA should point to one of the standards in
the target list called 0000_throughput_standards.
ADC status (2007-March-7)
- For a description of the current status on the ADC please see:
ADC Current Status
LRIS Blue side shutter replaced (2006-Sept-7)
- The blue channel dewar is installed in LRIS following the
replacement of the blue channel shutter. The shutter that
was replaced failed due to ware between a pivot pin and arm
used to open and close the shutter. The new shutter was
tested and is operating normally.
As a result of the shutter replacement, the blue channel
dewar was realigned. As a result of the realignment, the FOV
is shifted slightly along rows and columns. This shift along
the horizontal axis for LRIS is roughly 70 pix. The shift is
benefits LRIS observers because it moves the bar closer to
the center of the gap between the blue channel
ccds. Observer's may notice that slits fell in the gap
during May to Aug because the dewar was slightly
The dewar is rotated 0.1-0.3 degrees such that the edges of
the imaging FOV are not parallel to the horizontal and
vertical array axis.
Grating station 3 is still unavailable. Observers will be
limited to three grating choices until the grating station
three is fixed. Thus, the fourth grating listed in an observer
configuration form will not be installed as it is assumed to
have the lowest priority.
LRIS Blue side shutter is stuck open (2006-Aug-29)
- The LRIS blue channel shutter has failed again, and work to
replace the shutter began on 28 Aug 2006. The LRIS blue
channel should be available in time for the next run which
starts 15 Sept. 2006. The shutter that recently failed was a
rebuilt shutter. We are replacing it with a new shutter. We
are replacing the shutter this week, and the blue channel
dewar will be re-aligned next week.
Observer who used LRIS during 21-27 Aug. 2007 should
carefully examine blue channel data to ensure the shutter
was not stuck open during their observations.
Blue channel FOV is slightly misaligned (2006-Aug-18)
- The dichroics used to redirect light toward the blue
channel are slightly misaligned. The misalignment results
in all slits being shifted roughly six arcseconds to the
right as seen on the CCD mosaic. This means that some
portion of a slit very near the slitmask bar will fall in
the gap between the two Blue channel CCDs. If this shift
significantly disrupts the primary science with your
slitmasks for the run starting 21 August, you may redesign and
submit a slitmask for milling.
Grating station three Fixed (2006-Aug-10)
- Grating station 3 is back in action. The encoder used to
help define the position of the grating tilt failed. The
failure mode was that the encoder could not find home. The
encoder was replaced with a spare. Observers should find that
the wavelength offsets for grating station three that were
measured during past runs may no longer apply. Observers may
again select up to four different gratings for their run.
Grating station three failed (2006-Jul-18)
- Grating station 3 is temporarily out of service. The
grating station failed 18 July, and has been taken out of
service. Observing teams are limitted to 3 gratings until
the grating station is fixed.
Dewar Service and Focus Change (2006-Jul-18)
- The LRIS dewar was removed from the instrument so that
summit staff could install a new dewar window, replacing the
original which was broken during the shutter repair in May. The
new window appears to work as desired. Following the work,
we've been adjusting the "tilt" position of the LRIS blue dewar
to improve the uniformity of focus across the field of view.
Currently the unformity in the blue CCD Y direction (along CCD
columns) is good, but the CCD X direction (along rows) shows a
gradient of about 100 microns from side to side. We judge this
to be an insignificant amount of defocus. Observers should be
aware that the nominal blue-side focus value has changed
significantly and is now approximately -2450. This should be
used as the central focus value when acquiring focus images with
the focus loop script.
New Web Pages Released (2006-May-15)
- There is a new look and feel to the LRIS instrument web pages.
The new web pages include information on the LRIS blue channel that
was previously lacking with the old pages. Procdures were updated,
conflicts in the information or procedures were resolved, and the
page information was re-organized. The LRIS documatation is grouped
into three primary stages: Pre-Observing when you will
require information for preparing proposals and materials in advance
of an observing run; Observing when you will be at the telescope, and
Post-Observing when you will need to process and analyze data in preparation
- The old web documenttion is obsolete and is no longer available.
Please report any errors, missing information, clarification requests,
and any problems with the pages to the lris instrument master.
LRIS Blue side shutter is stuck open! (2006-Apr-25)
- The LRIS blue side shutter has failed and work to replace
the shutter will begin on 5 May 2006. The shutter is stuck
open which permits observing. Several new scripts are
available to observers for use during afternoon setup and
science acquisition. The trapdoor is currently being used as
a shutter for the blue side. Red side data may still be acquired
simultaneously. Your SA will review the scripts with you
upon your arrival, but please see
trapdoor_as_shutter.html for a brief description of
the observing scripts.
LRIS Configuration Form Updated (2005-Nov-16)
- A new version of the LRIS configuration form is available at
The new form is customized to the number
of nights in your observing run, links to the slitmask database,
and populates dropdown lists with facility and user masks. For LRIS,
the blue side elements are listed, and a default configuration is
presented. When prepairing for upcoming LRIS and DEIMOS runs,
please use the new configuration forms.
IRAF script updates (2005-May-19)
- A new version of lrisbias.cl is available in the
to handle the 4-amp case (old version only worked with 1 or 2 amps).
The script was tested with both red and blue images. Please report
any problems with the script to the LRIS master. Chris Fassnacht
is the author of the new version of lrisbias.cl.
IRAF Changes (2004-Oct-12)
- LRIS accounts have now been upgraded to use v2.12. Also,
the tasks previously found in the ucsclris package
have now been moved to the keck.lris package. Simply
type keck and then lris to load the new
July 28, 2004 New telescope focus procedures (MIRA)
- A new MIRA script will be released for the August 2004
run. The script, which allows more flexibility, error checking and
greater efficiency on sky, will be run by the OA or by the SA at
least once per night. In addition, one of the MIRA runs of the
night, preferably the first, will involve taking focus mode images
on red and blue side simultaneously and running a double side
focus mode with before and after stacked images. The MIRA analysis
will be performed using the images of one side only, red if the scientific program is
exclusively using the red side, blue
if the scientific program uses the blue arm or both
arms. The extra time required to run this script is gained back by
the ability to configure LRIS while the telescope is
slewing to the target star. Subsequent MIRA runs will be done with
the minimal amount of reconfiguration and number of
An additional advantage of the new MIRA script is the possibility for
the astronomer to specify a user setup to reconfigure LRIS after the MIRA
run: the old MIRA script used to reconfigure LRIS as it was before running
the MIRA script, which in most cases was not ideal. In order to specify a
custom post-MIRA instrument setup follow the following steps:
- In the afternoon, after the UT date has changed, configure LRIS
as you need for your first post-MIRA observations (example for
slitmask alignment in the blue and spectroscopy in the red).
- In a punaluu xterm type setMira
- The script will generate the file
- After sky focussing, when you have the correct focus values for
the red and blue side, you can edit the file.
Do not hesitate to ask your SA in case of questions. The new MIRA
script and its features are being tested for the first time: we encourage
feedback from the users.
- July 13, 2004 New Instrument focus procedures:
- During the June run we have collected evidence of
focus variations during the night. We are investigating the problem and the
July 13 Engineering night will be partly devoted to the problem. In the
meantime, we strongly encourage observers to follow our recommendation for
- Dome focus: data will be collected by the support astronomer in the
late afternoon as close as possible to opening time.
- Sky focus: the support astronomer or the observer will collect a sky
focus sequence with both red and blue arms with dichroic 560, red
filter: R, blue filter: G. Different setups may be used if beneficial to
the science program. Please consult your support astronomer.
- Set the focus values determined on sky using the XLRIS GUI and
issue the following command on the punaluu window:
- focus_table > focus_yymmdd_RG560
- lpr -P lw4 focus_yymmdd_RG560
- Use the printed set of focus for your science setups.
- June 12, 2004 - The LRIS red side has been repaired and will be
available during the June run. Noise and conversion factor have been
determined for both amplifiers. Please check the new values on the
- June 8, 2004 - URGENT MESSAGE:
THE RED SIDE WILL PROBABLY NOT BE AVAILABLE FOR THE JUNE RUN.
- The red side electronics problems (see below) has recently
worsened. Despite our efforts to fix it, it is unlikely that June
observers will be able to use the red side.
Please contact the Instrument Scientist if you have questions.
- Red Side electronics problems (January 2004 - ongoing)
- Since January 2004 the red side electronic is not working properly. The
problem, dubbed " The red side saturation", which usually
manifests itself after a power-cycle of the instrument red side
electronics, generates completely saturate images (including pre and post
pixel areas). We have not yet found a way to fix the problem. We know that
it clears itself after a period of time. This period, which initially was
of 2-3 hours, has, since April, lengthened: it takes now more than 12
hours to get the red side back. We have been working to fix this problem
since it appeared, but till now we have failed to find a permanent cure.
The observers are advised that they may loose the red side capability
at any time, but especially in these situations:
1) First day of the run - When LRIS has to be power cycled for
installation on the telescope.
2) Every time a red electronics power-cycle is needed.
- As a temporary measure, till
the problem is fixed, we will anticipate LRIS installation at the
telescope as early as possible and we have advised all personnel to limit
the chances of events by minimizing the
power-cycling of the electronics. Astronomers are requested NOT to
attempt any power-cycling without talking to their SA.
- Blue camera oil leak (May 2004)
- During the May 2004 run a new oil leak in the blue camera was
discovered. The effects of the leak (partly smeared images) could
only be seen during MIRA acquisition, when the telescope in PMFM
mode. The leak from the back of the first triplet affected only the edge
of the lenses. No effects were noticed in the images taken at PMFM=0.
During the on-deck period at the beginning of June, the camera was taken
off the instrument, disassembled and cleaned. The leaking seal has also
been replaced. On June 4 the blue camera was back on-board.
- New LRISB pointing origin (19 November 2003)
- A new pointing origin named LRISB has been defined. It is meant for
users who want to do imaging in the blue side. The LRIS pointing
origin can is still available and puts the object at the center of the red
side detector, which falls in the gap on the blue side. This LRISB pointing
origin puts the object at approximate (x,y) coordinates as listed in
the table below
- Changes to the XFOCUS widget (31 Oct. 2003)
- Minor changes to the XFOCUS GUI and procedures:
- 1) The GUI can still be started from the background menu from the
Xfocus menu: this menu now has two entries, one for the red side and
one for the blue side. The GUI will automatically set the correct
number of holes to be looked for in the focus images (44 in the blue,
68 on the red side).
2) The focus procedure
now uses the correct pixel scale for the blue side (0.135"). This
result in a smaller average FWHM for a well focused point (~0.23)
- Blue images shifted on the blue side (Sep-Oct. 2003)
- During the late September run a horizontal shift (right-to-left) of
approximately 135 pixels was noticed on the blue CCD. The problem affected
only the MOS observers, who had slits close to the right hand side of
the gap. The problem was due to a camera misalignment The blue camera had
been taken out to repair the blue shutter which had broken in August and
to fix a couplant-oil leak . The installation of the camera is a difficult
procedure which may be prone to errors. The camera was re-aligned in time
for he start of the late October run.
- Adjustment to Pickoff Mirror Coordinates (9 Oct 2003)
- Observers who design slitmasks with pickoff mirrors using
the UCSC software package need to be aware of a change. This
software provides the pixel location onto which you should
place the alignment star in order to get your objects into the
slits. During a recent engineering night we re-measured the
offset between the predicted and actual location of stars on
the guider. Until further notice, observers should apply
the following offset to their pickoff star coordinates:
These represent only a small change from the last measurement
made 3 years ago.
Changes to LRIS Slitmask Submission Process and
- The procedures and deadlines for submitting LRIS slitmasks
are changing effective 2003-Oct-04 due to the onset of
DEIMOS slitmask milling at the Keck summit. Key points:
for complete information on the changes.
- The existing FTP-based submission system is being
eliminated in favor of a web-based slitmask database system
which offers numerous advantages for both observers and the
observatory. Observers will need to register with the
database in order to submit slitmasks.
- Due to the heavy increase in slitmask milling workload
anticipated with the inclusion of DEIMOS masks, the deadline
for submitting both DEIMOS and LRIS slitmasks will increase
from 3 to 4 weeks, effective for all LRIS and DEIMOS
observing runs starting in October 2003. Particularly
complex masks (200 or more slits) must be submitted 8 weeks
The change will affect all LRIS users whose run starts after
Oct. 20 2003.
LRIS-R saddlebag electronics repaired (Sep 15,
- The red side electronic failed during some test while LRIS
was on the deck. It turned out that one of the amplifiers on the
analog board was the cause of the malfunction. The electronic
has been repaired. We believe that the gain, the saturation
level, the bias levels and CTE have been restored to their
previous values. New measurements will be done ASAP.
LRIS-B shutter REPAIRED (September 13, 2003)
- The blue shutter broke again during the August-September
run. It has since been repaired using available spare
parts. The shutter is now working nominally.
Status of the blue side mechanisms (August 2003)
- The LRIS servicing
mission has started. The blue side mechanisms have been inspected and
- Grism: the grism transport mechanism suffers
from vibration and bar-stooling problems. A temporary
solution has been implemented and will allow normal
operations of the mechanism in the near future. A complete
repair mission is being prepared.
- Dichroic and Filter have been partially
serviced. No major problems have been found. In June the
dichroic latch mechanism caused some sky-time loss. The
filter failure during the engineering night of July 2 was
caused by a loose screw trapped between a filter cell and
its retainer. Both mechanisms have been repaired.
- LRIS red side detector electronics problems - July 2003 - During
the engineering night of July 2nd we collected data to generate transfer
curves for the CCDs. The red side curve showed that blooming effects were
affecting the images as soon as the ~34,000 ADU limit was passed. A repair
mission was initiated immediately and a temporary patch allowed
observations with close to nominal performance for the CCD at the end of
July. During the August on-deck time the repairs where completed and as of
August 7th the red CCD is working nominally. Refer to the
page for further details.
- LRIS-B shutter REPAIRED - January 22, 2003
The broken blue shutter was replaced at the end of the
January run with the available spare. The shutter is now
LRIS-B shutter failure
- Currently (Dec. 29) the blue shutter in LRIS has failed. It is partially
open, and we may be able to force it completely open. It is not expected to
be fully repaired until after the current LRIS run ends. In the meantime,
observers may want to consider changes to their observing programs.
A Web page describing some
workarounds and extra calibrations can be consulted for ideas.
- The script PMlog allows to produce an observing log in
electronic format. It can be called by any lris account
from any machine (either local or punaluu).
The script is very simple and produces either ths standard
logsheet information (taken from the fits headers) or any
custom combination of keywords. The print option generates
a postscript file which can be printed with the lpr
command. Users are encouraged to use it and give feedback.
Usage : PMlog directory [-p filename] [KEY1] [KEY2] ... [KEYn]
Examples: PMlog . RA DEC Ut TARGNAME object
PMlog /s/sdata223/lris2/9Sep2002 -p
filename RA DEC Ut TARGNAME object
Options : -p filename prints to file filename.ps in the data
- With the NBC commissioning the focus tables for LRIS,
both red and blue side, have changed. To facilitate the
users, the new script "focus_table" has been made available on
The script calculates the focus table for the night, for each instrument
configuration. It assumes that the correct focus was
measured for the current instrument setup (by means of the
usual focus loop procedure) and it calculates the focus
for all the settings applying a list of recently
calibrated offsets. It does NOT correct for the difference
between dome focus ad sky focus (-20 microns, to be added
to dome focus values).
LRIS update for september
2002 (18 Sep 2002)
LRIS update for august
2002 (15 Aug 2002)
IRAF mask alignment software updated
- To support slitmask alignment using the New Blue Camera,
Alice Shapely and Chuck Steidel of Caltech have generously
provided CARA with an updated version of Drew Phillips's
UCSCLRIS package for IRAF. The revised version is installed
on the headquarters IRAF account; observers will get this new
version (2.0) when they load UCSCLRIS. The package features
three new tasks:
These tasks account for the spatial gap between the two CCDs
in the mosaic. Also, the default imtool buffer size for LRIS
observers has been updated to 4096×4096 to support the new
mosaic size. Please advise us of any problems you encounter
in running this new software.
- xboxb performs slitmask alignment on blue
- mshiftb performs offsets in pixel space on
the blue mosaic
- check_boxesb is a front-end or
xboxb allowing the observer to define or
preview alignment box locations
Changes with the LRIS New Blue Camera (NBC)
Efficiency curves with the LRIS New Blue Camera (NBC)
Gain and Focus measurements with the LRIS New Blue Camera (NBC)
Wavelength coverage and throughput with the LRIS New Blue Camera (NBC)
Efficiencies of LRIS 600-line Gratings
- Plots and data showing the relative efficiencies of the
three 600-line gratings available with LRIS are now
available on the Red Gratings
LRIS Update for
February 2002 (21 Feb 2002)
LRIS Update for
January 2002 (24 Jan 2002)
Enjoy Faster MALIGN (06 Jan 2002)
- Changes to the data gathering component of MALIGN have
sped up the procedure by about 1.5 minutes for each
single-sided alignment and nearly 3 minutes for
double-sided. The modifications are:
Note that this requires that the telescope execute a
small (30 arcsec) offset at the beginning and end of the
procedure in order to center the image on the left
- only half as many CCD rows are read out
- images are read using dual-amp readout
LRIS Update for
December 2001 (13 Dec 2001)
Movable Guider Repaired (26 Nov 2001)
- The offset guider, which is used for guiding during
imaging and multislit spectroscopy, failed during the November
LRIS run due to a mechanical breakdown. The guider has now
been repaired and is once again in full working order.
Observers should bear the following in mind:
- Despite an active maintenance program, LRIS is a
complex instrument with many moving stages which often
experience mechanical problems.
- You can minimize the risk of losing time during
observing runs by coming prepared with backup plans for
imaging and longslit spectroscopy.
- Designing your slitmasks to use pickoff mirrors will
help ensure that the loss of either guider does not
impact your multislit observing program.
New Blue Camera Delayed (26 Nov 2001)
- The deployment of the new UV-hot detector and dewar for
the blue side of LRIS has been delayed. The target date for
commissioning the system had been late November 2001, but
recent technical problems have made it impossible to achieve
that goal. Although a new schedule is not yet available, the
system will not be available in December and perhaps not for
some months afterward.
LRIS Update for
November 2001 (01 Nov 2001)
New Tool Allows Observer Control of Dome Lamps
(31 Oct 2001)
- By using the new Dome Lamps tool developed by Bob
Goodrich, observers can now turn the spectral and imaging
flatfield dome lamps on or off remotely. The new GUI is
available by clicking on the CMD...
button on XLRIS and selecting the Control
dome lamps option. This capability is intended to allow
Observing Assistants (OAs) to leave the mountaintop as
scheduled at the end of the night while still enabling
observers to obtain calibrations. Please note the following:
A corresponding domelamps
command allows control of the dome lamps from observer
- Assistance from the OA is still required to place the
telescope at the dome flat position
- Observers must complete all calibrations by 8 a.m.,
at which time the day crew will begin to arrive.
LRIS Collimator Mirror Cleaned After Spill (25 Oct 2001)
- Recent reports of vignetting of the LRIS imaging field
were traced to a fluid spill on the LRIS collimator mirror.
The mirror was removed from the instrument and successfully
cleaned by summit staff without the loss of additional time,
and performance of the mirror should once again be normal. It
seems likely that the spilled fluid may be optical coupling
fluid leaked from the LRIS blue camera; hence, we are
investigating whether any changes in image quality are evident
on the blue side of the instrument.
Slitmask Alignment Checklist Available (19 Oct 2001)
- A new checklist
which comprehensively describes the steps required to align a
slitmask is now on-line. The procedure includes some
operations to shave valuable time off the alignment process,
and has been observer-tested. This checklist is designed for
the most common slitmask alignment technique, which is:
- alignment boxes included on slitmask;
- no pickoff mirrors in use (coarse alignment via
- using xbox software for
Dewar Warming/Cooling Curves Available (1 Oct 2001)
- Plots of the dewar
temperature during warmup and cooldown
phases have been posted online.
Ghost update (18 Sep 2001)
- Chuck Steidel recently studied the ghosting problem on
the blue side and reports some new findings: ``The brightest
'ghosts' on the blue side are not reflections from
the back side of the dichroic as we previously thought. They
are actually dispersed light from the red side (if the flat
mirror is in on the red side then you get only single ghost
images, if a grating is in then the ghost spectrum moves by
the same amount as it does on the red side in responding to a
change in the grating angle). The dispersion in the primary
ghost spectrum suggests that it is second-order dispersed
red-side light. All of this suggests that we may be able to
improve this situation by improving the baffling internal to
the instrument. Incidentally, the ghosts also depend on
whether a grism is deployed on the blue side, suggesting that
the reflections that are making it into the blue camera are
'channeled' there by the grism.''
Safeguard Added to 'xbox' (12 Sep 2001)
- The xbox task used for LRIS
slitmask alignment has been changed to account for the
inability of the guider to operate properly when rotating the
instrument. When you run xbox,
it will now prompt you to have the OA disable guiding during
the move, and re-enable it after the move completes. You must
acknowledge that guiding has been turned on and off by
pressing the <Enter> key; these additional
prompts may cause longtime LRIS observers to believe that
their moves have been sent when, in fact, the program is
waiting for input. Please pay close attention to the messages
you receive from xbox.
New Tool Aids Slitmask Alignment (11 Sep 2001)
- Observers who use the xbox
program for slitmask alignment will benefit from the new task
check_boxes which provides an
interface for defining and inspecting the pixel locations of
the alignment boxes. The task allows you to generate a
coordinate list rapidly and to see immediately whether the
coordinates you gave agree with the box locations.
LRIS Low-level Software Repaired (08 Sep 2001)
- Diligent work by Keck software staff has tracked down a
bug which prevented the low-level software from functioning
appropriately. Keywords should no longer lie, and XLRIS
should now give accurate information about the state of the
instrument. We will be disabling the monitor tasks which were
needed to deal with the bugs, and which caused numerous
spurious warnings which confused observers.
LRIS Update for
September 2001 (06 Sep 2001)
Select Guider Filters Yourself (31 Aug 2001)
- A new GUI allows observers to show or change the selected
filter in each of the LRIS guider filter wheels. The GUI can
be invoked from the CMD... menu on
the XLRIS GUI; click on the item labeled Select guider filter.
LRIS Update for July
2001 (26 Jul 2001)
LRIS Update for May
2001 (17 May 2001)
Software Changes (26 Apr 2001)
- Changes made to the LRIS software in early April are
believed to have been responsible for some abnormal behavior
of the instrument, including:
We have undone the change to the keyword library software
which appears to have been responsible for the problems;
however, this has had a side effect on XLRIS that prevents
it from reporting clear when the dichroic, grism,
or blue filter has been removed from the beam. Instead, the
GUI will report the name of the currently selected element
whether the elements is in the carousel or in the beam. To
determine the position of the blue elements, you can bring
up a window which reports the position of the transport
mechanisms by selecting the item LRIS Transport
Status on the LRIS Eavesdropping menu
from the OpenWindows pulldown menu.
- XLRIS window freezing up every time a grating is
- watch_ccd occasionally freezing,
preventing frame numbers from incrementing and causing
loss of data
LRIS Update for April
2001 (19 Apr 2001)
A Better Way to Run Xfocusloop (13 Apr 2001)
- It is no longer necessary to type the
xfocusloop or xfocusloop.blue commands
to acquire internal focus images for analysis with the
Xfocus widget. Observers can now start
a focus sequence on the red or blue side by going to the
appropriate Xpose window and
selecting the Focus loop option on the
CMD... menu. Also, please note that red and blue
focusloops can be run simultaneously to save time.
New Arc Lamps Aid Blue Wavelength Calibration
(12 Apr 2001)
- Arc lamps containing zinc (Zn) and cadmium (Cd) have
replaced the previously-used krypton (Kr) and xenon (Xe) bulbs
in the LRIS calibration lamp complement. The new lamps should
provide sufficient wavelength coverage to enable accurate
calibration of all blue spectra. Observers must bear in mind
that these UV-emitting lamps take time to warm up: the Zn lamp
reaches half-peak output in 150 sec, and the Cd lamp requires
250 sec to reach the half-peak-flux level. Neither lamp
emits significant levels of UV flux in the first minute of
G Filter Now Available (12 Apr 2001)
- A new G-band filter was recently received and
has been installed as a permanent addition to the blue side of
the instrument. This filter has a transmission curve which is
broader than the Gunn g' filter but has no color term
relative to g'. See
& Hamilton (1992, ApJ, 394, 81) for further discussion
of this filter.
LRIS Software Changes (12 Apr 2001)
- We took the opportunity during the last two weeks to
complete several changes to the LRIS software. Changes of
interest to observers include:
- XLRIS blue controls simplified. The
dichroic, grism, and blue filter stages previously had
separate controls for the transport and carousel. These
have been replaced by the single control which indicates
clear when the filter, grating, or dichroic is
stowed and out of the beam. Also, the blue focus window
slider bar now corresponds to the legal range of blue
- Keyword behavior changed. All queries
for instrument keywords will now cause an enquiry to the
corresponding instrument stage instead of looking up the
value in a state table. This should fix most of the
instances of "lying keywords" recently encountered.
- Blue optics now clamp reliably. The clamps
which hold the blue optics in position in the blue beam
have not always fired reliably on the first try. The
motor control software has now been upgraded to check
the status of the clamps and apply them up to 5 times as
needed to ensure that the clamp is activated. This
change should eliminate the problems seen previously with
undefined blue optics.
New Warnings for Keywords and Clamps (21 Mar 2001)
- LRIS is currently experiencing some problems with
keywords and blue mechanisms:
- Lying keywords. Sometimes the LRIS keywords
such as slitname (name of the current slitmask)
or graname (name of the current grating) fail
to update when the corresponding stage is moved. This
can lead to confusion, because XLRIS will then display
the wrong name for the current slitmask, filter,
grating, etc. Furthermore, the keywords written to the
image header will be wrong.
As a stopgap measure, the LRIS software has been
revised to check every 60 seconds for invalid keywords,
and to warn observers via the TkLogger window
when invalid keywords are found. Observers can then fix
the problem by executing the fixkeywords command
from a punaluu xterm window. We will
investigate this problem more fully with help from
Caltech software support in April.
- Undefined blue optics. The optical
mechanisms on the blue side (dichroics, gratings, and blue
filters) are held in position by pneumatically-driven
clamps. Recently we have seen evidence that these clamps
do not always activate, leaving the optics undefined
mispositioned. This can produce shifting
of the resulting images or spectra on the blue side, or on
the red side if a dichroic is in use.
We will update the motor control software to deal
with this issue as soon as possible. Meanwhile, the
LRIS software will check every 60 seconds for
unclamped blue optics and will warn observers via the
TkLogger window when any are found.
Observers can then execute the check_clamps
script from a punaluu window to activate the clamps.
LRIS Collimator Recoated (16 Mar 2001)
- The collimator mirror at the heart of LRIS has received a
new high-tech coating and been reinstalled into the
instrument. Prior to the LRIS-B upgrade the collimator was
coated with silver, but this was changed to aluminum in
mid-2000 in order to provide better UV reflectance. During
the two-week LRIS downtime in March, this mirror was coated
with a hybrid silver-on-aluminum coating at Lawrence Livermore
National Laboratory. The new coating melds the best of both
worlds, providing better UV reflectance than Al and red
response comparable to silver. The chief effects observers
should see are:
We hope to redetermine the mapping from pixel space to the
sky in the near future.
- Improved throughput at all wavelengths.
- Possible changes in the plate scale and astrometric
mapping of the LRIS field-of-view due to realignment of
LRIS Guider Filters Changed (16 Mar 2001)
- A new set of broadband filters is now installed in both
the slit-viewing and offset guiders. The set includes clear,
B, V, and R filters. See the guider
filter page for details.
LRIS User Script Library (12 Mar 2001)
- LRIS now has a complete
library of commands which can be used to control the
instrument and telescope from the command line as well as from
user shell scripts in order to perform complex operations.
For more information, see LRIS Shell Scripts.
LRIS Update for March
2001 (08 Mar 2001)
Blue Side Filter Changes (26 Feb 2001)
- The V filter for LRIS-B has been temporarily
returned from the manufacturer in order for us to verify that
the filter delivers double images. It has been re-installed
into LRIS but is not expected to provide scientifically
useful imaging capability. We expect to return it to the
vendor shortly. The blue
filter list has been updated accordingly.
New LRIS-R&B Logsheet (26 Feb 2001)
- A new logsheet suitable for
simultaneous red- and blue-side operation is now available
online. It and others can be accessed from the
LRIS logsheets list.
Dichroic Data Posted (21 Feb 2001)
- A list of LRIS dichroics is
now available, featuring transmission data for the four LRIS
dichroics installed as part of the blue-side upgrade.
LRIS Update for February
2001 (07 Feb 2001)
LRIS FAQ Available (30 Jan 2001)
- Answers to some frequently asked
questions have now been assembled on-line for your
convenience. While still embryonic, this version may help
shed light on certain issues related to LRIS-B.
Grism Data Posted (30 Jan 2001)
- A list of LRIS-B grisms is
now available, featuring wavelength coverage and efficiency
data for the four LRIS-B grisms.
Blue Side Filter Changes (26 Jan 2001)
- The blue side of LRIS has lost one filter and gained
another. The V filter, which has been rendered
scientifically useless by delamination of the filter sandwich,
has been returned to the manufacturer for servicing and is
expected to be gone for several months. However, Chuck
Steidel has generously supplied his Palomar U filter
and it is now available for use. Prospective observers should
note that this U filter is undersized and will
significantly vignette the beam, passing only about 57% of the
light. Chuck also reports that the G filter he has
ordered for public use on LRIS-B (4730/1100) is expected to
arrive in time for use during the March LRIS run.
Single-sided MALIGN Working (17 Jan 2001)
- MALIGN-ing the telescope using only one set of PMFM
exposures (a.k.a. ``single-sided MALIGN'') should work again
now that we have redetermined the appropriate correction
factor, and will become the default mode for running MALIGN.
Completing a single-sided MALIGN takes only about half as long
as the 2-sided MALIGN which we have been running lately on
LRIS. If the primary mirror stack is suspect, two-sided
MALIGN should still be run. Observers who prefer two-sided
MALIGN for reasons of security are still welcome to request
that the OA use it.
LRIS-B Overview Available (15 Jan 2001)
- An LRIS-B Overview is now
available thanks to Chuck Steidel, LRIS-B P.I. It provides
a succinct introduction to the characteristics and vagaries
of the new blue side of the spectrograph.
Don't Rotate While Guiding (14 Jan 2001)
- A simple test tonight confirmed what some slitmask
observers have seen for some time: if you attempt to rotate
the instrument while guiding, the guide box will not follow
the object. Until further notice, follow these steps when
We are investigating the reasons for the failure.
- take your mask exposure
- determine the solution (e.g., run xbox but do
NOT send moves if a rotation is suggested
- ask the OA to turn guiding OFF
- send the move involving rotation
- ask the OA to resume guiding
- check alignment again
XLRIS-Generated Scripts OK (14 Jan 2001)
- We believe that scripts generated using the
WRITE button on XLRIS now work properly; i.e.,
they do not arbitrarily insert blue optical elements into
the optical path.
LRIS Update for January
2001 (11 Jan 2001)
XLRIS Bugs (03 Jan 2001)
- Observers should be aware that LRIS currently has two
bugs in it that can affect observing:
Update on Secondary Misalignment and Optical
(24 Dec 2000)
- Further testing of optical alignment suggests that the
secondary is not significantly decentered from LRIS. In
fact, we suspect that the rotator compensations to secondary
piston were incorrect, resulting in significant focus change
of LRIS with rotator angle. The secondary piston
corrections have now been disabled. Observers are advised
to monitor guider focus frequently in order to determine
whether focus has degraded since the last MALIGN.
Also, please be advised that single-sided MALIGN is
currently not returning correct values for secondary
piston; hence, only double-sided MALIGN is recommended for
use on LRIS until further notice. (note: fixed as of
15 Jan 2001)
Observers have recently reported that LRIS images
exhibit significant distortion, even after running
MALIGN. The effect is most noticeable on the leftmost
part of the images. If your program consists primarily
of direct imaging, you should ask the OA to disable
LRIS tilt compensations to obtain the best
image quality. Note that with tilt compensations
disabled, focus will be expected to degrade more quickly
than normal as the LRIS physical rotator angle
(ROTPPOSN) changes. Check focus carefully and be
prepared to re-run MALIGN when the physical rotator
angle changes by 45° or more.
Blue Side Unavailable for December LRIS Run
(24 Dec 2000)
- The LRIS blue dewar has a faulty limit switch on its
focus stage, preventing the dewar from focusing properly.
Unfortunately, removal of the red dewar is necessary in
order to make the repair. Since we do not want to put the
red side at any risk for this run, we will not repair the
switch until after the end of the current LRIS run on 1
Jan 2001. We hope to complete the repair before the
mid-January LRIS run begins.
New Utility for Writing Tapes (18 Dec 2000)
- A new program called mktartape makes writing tapes of
your data virtually foolproof. The program has many features
designed to make the process easier and prevent observer
mistakes, such as:
The script can be run from any observing account by going to
the Openwindows pulldown menu, selecting the submenu called
Instrument Utilities, and clicking the item
labeled Make tar tape.
- verifies the existence of the data directories entered by user
- allows observer to select any tape drive on the current
machine without needed to know the name of the tape drive
- allows user to append new files to those already on tape
- can write multiple data directory trees to tape in one pass,
putting each into a separate tarfile
- after writing data to tape can (optionally) verify tape
while creating a tape log
- optionally rewinds and ejects tape at completion of writing
LRIS Update for December
2000 (14 Dec 2000)
Keck I Shutter Fixed -- Restrictions Lifted (7 Dec
- The Keck I dome shutter is now repaired and observers can
once again gain unvignetted access to the zenith.
Policy on the use of LRIS-B during Semester
2000B (1 Dec 2000)
LRIS-B has essentially completed its commissioning phase and
is expected to be ready for regular use by observers
beginning with the 2001A semester in February 2001.
However, the instrument is not yet in its final state:
Furthermore, the instrument documentation has not yet been
received and the Instrument Specialists are not all trained
in the use of the blue side. However, the instrument should
be simple to use and can prove quite useful to observers for
slitmask alignment and other applications. Thus, our
current policy on the use of the blue side of LRIS is that
the instrument is available only on a SHARED RISK basis. In
practice, this means that observers are allowed to operate
the blue side mechanisms and detector, but that Keck cannot
fully support the instrument. So if (for example) the
detector system hangs or a blue-side mechanism fails, we
will not take extraordinary measures to fix them during the
night. Of course, once the instrument is documented and all
personnel have been trained, we will begin supporting the
instrument fully. We hope to institute full support by the
end of Semester 2001A (end of July 2001) as scheduled.
However, we can offer no assurance as to when any of the
hardware problems -- collimator coating, U filter, or
detector -- will be resolved.
- The collimator remains coated with Al; we hope to
increase red throughput by applying a silver-on-aluminum
coating in March 2001. Instrument throughput is expected
to change at that time, increasing by 10% in the red.
- The need to remove the collimator for recoating means
that the plate scale (for both the blue and red sides)
could potentially change.
- The U filter has delaminated and been returned to the
factory for repair with an unknown return date, leaving
the blue side with only B and V filters at present.
- There is as yet no science-grade detector for the
blue side. The current detector is a very good
engineering-grade SITe 2048×2048 chip very similar
to the long-standing red side detector but with a little
bit more blue sensitivity. We do not know when a
UV-optimized science grade detector will be available.
LRIS-B U Filter Out of Order (1 Dec 2000)
- The U filter delivered with LRIS-B was found during
commissioning to have severe optical problems. The
manufacturer evaluated the filter and determined that it had
delaminated. A new filter is begin fabricated, but will
require custom optics to be made in Germany. A realistic date
for receiving a replacement is June 2001.
Rs Imaging Filter Available (27 Nov 2000)
- A new public "Rs" filter is available for LRIS observers
thanks to the generosity of Chuck Steidel (Caltech). This
broadband filter features higher peak transmission than the
standard LRIS R filter (96% vs. 78%) and a sharper
transmission profile. Although inspired by the so-called
"Spinrad Night Sky Filter" in use for many years at Lick, this
particular filter (generally referred to as the "script R" or
filter in the
literature) was originally defined by Mark Dickinson and Chuck
Steidel in order to cut off the substantial red tail inherent
to the Spinrad filter. The Rs passband is intended to be a
nice compromise between R and I while maintaining a dark sky
background (21.0 AB mag/arcsec^2 under moonless conditions at
most sites, including those with broad Na D problems like Lick
and Palomar). Its transformation to standard systems is given
and Hamilton (1993). Thanks to Grant Hill, plots and
transmission curves are accessible from the LRIS General Filter List.
Run Copymon Automatically with "Autocopy" (22
- Bob Goodrich has kindly automated the process of starting
the copymon routine for mirroring
your LRIS images to a HQ disk. If you select the Autocopy option under the LRIS Widgets item on the OpenWindows
menu, it will select the best scratch disk, start copymon writing to that disk, and
optionally copy over existing images (copymon only grabs
Secondary Misalignment and Telescope Focus (21
- All observers should be aware that the K1 f/15 secondary
remains out of adjustment. We discovered last month that the
cams which support the secondary were loose, and hence the
position of the secondary has been changed by an unknown
amount and must be recentered. We've acquired some
engineering data which should allow us to measure the severity
of the effect, but in the meantime you must assume that the
secondary is off center relative to the instrument rotator.
This is particularly important to bear in mind for LRIS
observing, because it means that the rotator-dependent
correction factor to the telescope focus is likely to be
inaccurate. Until further notice, observers are advised to
run autofoc or MALIGN whenever the PHYSICAL rotator position
changes by 30 degrees or more since the last calibration.
Arc Lamps and Movable Guider (20 Nov 2000)
- LRIS observers who calibrate their data using the
internal arc lamps should be aware that their wavelength
solutions may depend on the position of the movable guider
stage. The arc lamps are physically mounted on the same stage
which holds the movable guider; hence, as the movable guider
changes position, the line lamps move in the direction
perpendicular to the slits (i.e., in the wavelength
direction). The result is that the illumination of the
slitmask by the arc lamps may depend on the movable guider
position (header keyword TV1FPOS), although we have
not yet quantified the effect. Observers who require accurate
wavelength solutions should:
- Ensure that the movable guider stage is always set to
the same position when taking arc lamp exposures. The
most uniform illumination is found when the guider is
near the low end of its range (TV1FPOS=216).
- Use night sky emission lines to verify wavelength
Slitmask Disposal (17 Nov 2000)
- Space limitations at the summit have forced us to change
our policy on keeping old LRIS slitmasks. Previously, we
saved all slitmasks in case they were needed in the future.
In practice, we've found that few of these masks were ever
re-used. As of October 2000, we will begin discarding LRIS
slitmasks immediately following a run unless the observer
requests otherwise. Note that we will permanently maintain
the design file used to manufacture the slitmask, allowing it
to be re-milled if it is needed in the future. Observers
wishing to have their slitmasks retained are welcome to
request for them to be mailed to their home institution
following the observing run.
LRIS Update for December
2000 (15 Dec 2000)
Keck I Dome Problems Limit Observations (6 Nov 2000)
- We recently experienced another mechanical failure with
the Keck 1 Dome Shutter, this time in one of the 8 cables that
moves the top shutter. Although troubleshooting has been
severely hampered by an extended bout of bad weather, we have
determined that it is safe to open and close the Keck 1 top
shutter, but with some restrictions on the range of movement.
Until the damaged cable is replaced, we cannot fully open the
top shutter. Movement is limited to a shutter position of 90
degrees, which means that the telescope beam starts to see the
top shutter at a telescope elevation of about 75 degrees.
Higher telescope elevations will of course see more of the
shutter, and the effects will of course be most severe in the
We have a replacement cable on rush order. IF the weather
cooperates, we hope to have the replacement cable
installed and properly tensioned in about 2 weeks.
Working on this problem requires opening the dome during
the day, so any bad weather will cause a direct delay in
the repair schedule. The shutter movement restrictions
will be in force until the repairs are complete and
tested. I will notify you again as soon as the repairs
Thanks for your cooperation and understanding during this problem.
Associate Director for Observing Support
New Version of UCSCLRIS IRAF Package Released (25 Oct 2000)
- Version 1.0 of the IRAF package UCSCLRIS has been
released for use. This package contains two tasks commonly
used by observers during the slitmask alignment process:
- xbox will check
the image header to determine whether images were taken
with the red or blue camera (INSTRUME keyword)
and will adjust the pixel scale accordingly to make
accurate offsets for slitmask alignment. Also, the safety
prompt will appear even if you run the task using the
:go directive in the parameter editor.
- mshift has been updated to prompt the user for
the camera name (red or blue) and to
set the pixel scale appropriately. (Note: the
movr and movb commands available
from the punaluu command line provide similar
capability and are recommended.) As with xbox,
the safety prompt will now appear even if you run the task
using the :go directive in the parameter
New LRIS Widgets Available (25 Oct 2000)
- Three new widgets are available from the LRIS OpenWindows
pulldown menu under the heading LRIS Widgets:
- Xfocus analyzes focus data
acquired with the focus_holes slitmask to
determine the internal LRIS focus.
- Xarcplots produces arc
line plots at any desired central wavelength and with
any combination of line lamps.
- Xdither allows observers to
define a series of spatial offsets and acquire images in a
pattern. This widget was previously available from the
LRIS IDL Tools menu.
LRIS Red & Blue Image Scales Measured (24 Oct 2000)
- LRIS-B P.I. Chuck Steidel (Caltech) has made measurements
of the new pixel scales on the red and blue sides. His
The scripts movr and movb which move
the telescope in red and blue pixel space, respectively,
have been updated with the new pixel scales.
- LRIS-R, w/dichroic: 0.2109 arcsec/pix
- LRIS-R, no dichroic: 0.2109 arcsec/pix
- LRIS-B, D560 dichroic 0.2151 arcsec/pix
Adjustment to Pickoff Mirror Coordinates (16 Oct 2000)
- Observers who design slitmasks with pickoff mirrors using
the UCSC software package need to be aware of a change. This
software provides the pixel location onto which you should
place the alignment star in order to get your objects into the
slits. Due to realignment of the slit-viewing guider, the
field of view has shifted slightly and hence, the coordinates
have changed. Until further notice, observers should apply
the following offset to their pickoff star coordinates:
LRIS Red Image Scale Changed (25 Sep 2000)
- Initial results suggest that the pixel scale of LRIS
images has changed by about 1.9% since the summer
refurbishment, which included adjustments to several elements
in the optical path. The new pixel scale appears to be around
0.211 arcsec/px, versus 0.215 arcsec/px prior to the upgrade.
This change has two important operational implications for
We have not yet derived a new plate solution for the
refurbished instrument, but will plan to do so after work on
recoating the collimator mirror is completed late in 2000.
We thank Judy Cohen for providing numbers on the pixel scale
- Because it relies on exact knowledge of the LRIS
imaging scale, the FIGDISP-based slitmask alignment
software will no longer work. Observers should plan to
use the xbox software package until we remap the
- Images acquired since the upgrade cannot be used with
the coordinates software to produce slitmask
designs. Observers must continue to use pre-upgrade
images in order to generate accurate coordinates for
Dual-Amp Readout Mode Restored (25 Sep 2000)
- Reprogramming of the Leach controller for the red side of
LRIS appears to have eliminated the previous problems with
freeze-up during dual-amp image readouts. Observers may now
return to using dual-amp readout mode for all LRIS images.
Observations Restricted to Single-Amp Readout
Mode (31 Aug 2000)
- The new LRIS CCD system has proven to be very prone to
getting stuck when operated in dual-amp readout mode. Although
procedure is well defined and the recovery time should be
short (5 minutes or less), the glitch results in losing part or
all of the image. Since single-amp-mode images do not show
this problem, observers are advised to use single-amp mode for
all nighttime exposures while we work to diagnose and fix the
'movr' Replaces 'imoffset xy xy' (28 Aug 2000)
- The FIGARO command imoffset xy xy, long used
by observers to move the telescope in LRIS red detector
coordinates, has been replaced by a new script call
movr. The syntax of the commands is essentially
the same, the arguments being:
Once the blue side is commissioned, a corresponding
movb command will enable moves in blue detector
- current x coordinate of target [pixels]
- current y coordinate of target [pixels]
- desired x coordinate of target [pixels]
- desired y coordinate of target [pixels]
LRIS-B Upgrade Status (14 Aug 2000)
- LRIS has been out of commission since early May for
installation of the new blue side of
the spectrograph. Work also included major changes to the
electronics and computers associated with the instrument. Due
to the diligent efforts of the LRIS-B team from Caltech and
many staff members at CARA, the installation of all optical
systems is complete and the red beam of LRIS is being prepared
to return to regular service as scheduled on August 22.
Long-time LRIS-R users will notice some minor
changes in the instrument, but red-side operations should
be mostly unchanged.
Although the LRIS-B optics are installed and working, the
science-grade CCD array for the blue side has been
returned to Caltech for servicing. It is still months
away from being installed and will also take some time to
commission and characterize. At this point, the
likelihood of releasing the blue side for science during
the 2000B semester is small; however, with luck we hope to
have the blue side available for shared-risk observing in
the 2001A semester.
See also the following items describing additional changes
which will impact LRIS users in 2000B.
New LRIS Internal Focus Procedures (14 Aug 2000)
- Recent testing during LRIS re-commissioning nights suggests
that the internal LRIS focus on the red side can vary
significantly (60 µm or more) on timescales of several
hours. Although the source of the variation is not yet
certain, it may be caused by temperature changes which affect
the red camera. It is likely that this has always been the
case, but we were previously unaware of this situation.
Furthermore, testing has shown that focus values obtained
from images with internal or dome illumination do not
correlate well with on-sky focus values; hence, we
recommend that focus be checked on-sky, ideally during
twilight. Until we have characterized the dependence of
focus on temperature, all observers should plan to
refocus the spectrograph during evening twilight and
possibly at various times throughout the night.
Since the focus values for each filter will not be
constant over time, we will no longer be posting
recommended focus values for each filter on the bulletin
board and the web pages; instead, we will post
focus offsets for each observing
mode and filter relative to the direct imaging R-band
focus. Observers will be expected to measure the R-band
focus and thus obtain correct focus in other modes based
on these offsets.
New software now allows the internal focus to be checked
in under 10 minutes during twilight illumination. However,
using this method requires that you have the
focus_holes.eng mask installed as part of your
slitmask complement; thus, you will have room for only 8
To summarize the new recommendations:
For further information, please
or consult the LRIS web pages.
- Observers will be responsible for measuring
- Plan to devote 10 minutes of twilight to checking the
R-band focus, as dome and internal illumination do not
adequately simulate the incoming beam from the sky.
- Request focus_holes.eng slitmask on your
LRIS configuration form to enable use of new focus software
Throughput decreased due to collimator
re-coating (14 Aug 2000)
- In order to provide acceptable UV throughput with LRIS
for using the blue side, we have replaced the old silver
coating on the LRIS collimator mirror with a new aluminum
coating. While this will help the blue/UV throughput of the
instrument, it also decreases the red throughput by about 10%;
hence, observers will require integrations lasting 10% longer
in order to achieve the same S/N in the red as with the old
LRIS collimator coating. By the beginning of the 2001A
semester, we hope to have the collimator re-coated with a
special silver-on-aluminum process which will provide
uniformly high throughput from the atmospheric cutoff in the
UV out past 1µm in the red.
Host Computer Upgraded (14 Aug 2000)
- The downtime for LRIS-B allowed us to upgrade the LRIS
host computer from an old SunOS workstation
(manuka) to a much faster new Solaris-based
workstation (punaluu). The following notable
changes to software were also made:
- XLRIS has changed significantly, with control
buttons now installed for blue-side mechanisms (dichroics,
grisms, blue filters, and blue focus); however, during
2000B observers should not move these mechanisms
unless given permission by the instrument specialist.
- Xpose has been upgraded but will work essentially the
same as before.
- FIGDISP has been upgraded for use with Solaris
and the old DataViews-based interface has been replaced
with Tcl/Tk controls. Note: the key
bindings for operations like FWHM measurements and row
plots have changed; please see the FIGDISP help page for details on
the new keybindings.
WMKOLRIS IRAF Package Release 1.0
(06 Jun 2000)
- Release 1.0 of the wmkolris
IRAF package is now available on the HQ network. The only
major change is an update to the lrisbias task which subtracts
overscan from images taken in either single- or dual-amp
readout modes. The task did not treat properly the case of
subwindowed data in dual-amp mode. Note that ESI observers
can also use this task to de-bias their data. Release 1.0 is
now available for downloading on the Keck FTP server.
LRIS Moves Back to Keck I (24 Feb 2000)
After a long sojourn on Keck II, LRIS has returned to
operation on the Keck I side, providing a true dark-time
instrument on both telescopes for the first time and
enabling eventual silvering of the primary mirror on Keck
II. The most was completed with only minor glitches
affecting startup on Keck I. Observers should bear in mind
that the western pointing limit required on Keck II
no longer applies, but that an eastern limit now
affects LRIS observations.
New Narrowband Filter Available (27 Jan 2000)
- A new public filter is available for LRIS observers
thanks to the generosity of Hy Spinrad (UCB). This narrowband
filter features a 100 Å-wide passband which
(unfortunately) varies radially, changing from a central
wavelength of 8536 Å at the filter center to
8605 Å at the outer edge. Plots and data
representing one cut across the filter are accessible from the
LRIS General Filter List.
Imaging Foci Recalibrated (02 Jan 2000)
During the LRIS engineering nights of 31-Dec-1999 and
1-Jan-2000, the focus values for all common imaging
filters were re-measured using night sky illumination.
The revised filter values have been posted online and incorporated into
the redfocus command to enable
automated focusing of the spectrograph from within shell
LRIS Red Shutter Accuracy Measured (17 June 1999)
- The finite length of time it takes LRIS to open and close
the shutter yields different effective exposures times at the
detector center verses the edges. Engineering test
T893: LRIS Shutter Timing Test was
recently completed in order to quantify the effect. This
systematic effect is less than 1% in 2 seconds with
LRIS. This does not necessarily mean that any single exposure
of 2 seconds is accurate to 1% in exposure time, because a
random component to the shutter opening time may
exist, which would increase the uncertainty.
Guider Filters Changed (16 June 1999)
- Since the LRIS slit-viewing guider filter wheel remains
inoperative pending software changes to take effect in late
June, the filter wheels for the slit-viewing and offset
(movable) guiders have been exchanged. The offset guider now
has available clear, B, R, and I filters, and the slit-viewing
guider is set to clear. The guider
filter list has been updated accordingly.
LRIS Scripting Macros Available (1 June 1999)
- Bob Goodrich has ported to LRIS many of the useful
macros which NIRC observers use to write their own observing
scripts. These commands
allow shell scripts to set up the spectrograph, define
exposure parameters, start exposures or exposure sequences,
and move the telescope. Please see the document
Script Control of LRIS for further
New Filter Transmission Curves (26 May 1999)
- Newly-measured transmission curves are now available on
the LRIS Filter List web
page. In addition to PostScript plots of the passbands, raw
data are available as ASCII text and FITS images. Also
listed are filter bandwidths and effective wavelengths.
600/10000 Grating Now Available (21 May 1999)
- A new gold-coated 600-line grating blazed at 1µm is
now available for general use with LRIS, courtesy of Alan
Stockton (U. Hawaii). Although we have not yet completely
characterized this 600/10000 grating, it is expected to
provide superior throughput to the 600/5000 and 600/7500
gratings in the near infrared.
LRIS Imaging Mirror Realigned (18 Mar 1999)
- For several weeks the LRIS imaging mirror has been out of
adjustment, causing the center of the imaging field-of-view to
lie well of the center of the detector. The mirror has been
re-aligned to put the center of the field of view (and the
slitmask crossbar) within 20 pixels of the CCD center. A
locknut was noted to be missing during this procedure,
possibly explaining the drift in angle.
ESO Spectrophotometry Catalog Link Available (16
- Esther Hu brought to our attention the
on-line catalog at ESO which gives finding charts and
spectral plots for various spectrophotometric
standards. The link is now available through the
Autoslit Documents Available in HTML Format (16
- The documents Introduction to
Autoslit and the Autoslit3
Manual are now available on-line in HTML format for your
WMKOLRIS IRAF Package "Alpha" Release
(16 Mar 1999)
- The first release of the wmkolris IRAF package is now
available on both the summit and HQ networks. The package
contains two tasks: lrisbias
subtracts overscan from images taken in either single- or
dual-amp readout modes, and lccdproc is a version of the
ccdproc task modified to use lrisbias for
overscan removal. Observers can access the package for data
reduction at Keck by starting IRAF and typing
wmkolris. The package is also available for
downloading and installation at home institutions.
New Tool Sets LRIS Red Camera Focus Appropriately
(15 Mar 1999)
- The redfocus command allows
scripts to always set the LRIS red camera focus to the
appropriate value, even when the focus values change. Given
the current filter and observing mode (direct, longslit, or
multislit) as command-line arguments, this script will read
the current focus values from a lookup table (maintained by the
LRIS Instrument Specialist) and set the focus accordingly.
The observer can also override certain focus settings by using
a ``personal'' version of the focus table.
Scripter Program Eases LRIS Setup-Script Composition
(15 Mar 1999)
web-based LRIS setup script composition tool has been released
for public consumption. This program will generate an LRIS
setup script based on the settings filled in by the observer
on the input form. All the observer needs to do is fill out
the form, download the script, and install it in the
~/setups directory on punaluu, the LRIS control
computer. The script can then be executed via the
GO button on XLRIS.
LRIS Focus Values Updated
(12 Mar 1999)
- Focus images taken during the last engineering run in
late February showed that the R-band imaging focus for LRIS
had changed by about 30 microns, a relatively minor change.
In response, the LRIS focus values web
page has been updated with all values increased by this
amount. More comprehensive data will be taken at the next
All Grating Stations Operational Again
(2 Mar 1999)
- All grating stations are now functioning normally again
thanks to the efforts of the capable summit techs. Grating
station 3, which had been flaky, was repaired and returned
to service. Observers may return to using 4 gratings
Movable Guider Slippage Fixed
(22 Jan 1999)
- The LRIS movable guider was repaired --- the problem
turned out to be a broken brake and a loose set-screw.
This should eliminate the slippage noted since December
and return operation of the guider to normal.
LRIS Now Initializes on Startup
(10 Dec 1998)
- The script to startup the LRIS control software has
been changed to include an optional LRIS initialization
step. Running this script when you first start the
instrument will return all LRIS settings to their default
configuration, undoing any changes that the previous
observer may have made. You may also wish to execute the
script on succeeding observing days to undo any changes made
by CARA staff during the day. The script window will appear
automatically once you select the "Start Up All LRIS
Control" option under the "LRIS Control Menu" pulldown menu,
and must be run or dismissed before other LRIS control
windows (Xpose, XLRIS, etc.) will appear.
LRIS Rotator Repaired
(07 Dec 1998)
- The badly degraded LRIS rotator bearing was replaced
during the first week of December and initial tests indicate
that the friction has been reduced to nominal levels.
Correspondingly, the maximum LRIS rotator slew rates have
been increased from 0.75 deg/sec to 1.5 deg/sec.
Intermittent Banding Noted in Spectral Flats
(30 Oct 1998)
Ratio of two flats
Spectroscopic-mode flatfields taken with the internal
halogen lamp have shown intermittent banding across the
illuminated portion of the chip. As shown in the
accompanying image, the symptom of this behavior is
diagonal banding seen in the ratio of two flatfields. The
banding is subtle, much lower in intensity than the
fringing already present in the chip, and thus may not be
visible in a single image. It can appear and disappear
within the span of a few minutes. Please report any
further instances of banding that you observe.
Get Notified When Readout Is Complete (14 Oct 1998)
LRIS users are familiar with the automated announcement
``exposure complete'' provided by the datataking
system. A new widget now allows you also to be
notified when the CCD readout is complete and a new
exposure can be started. The Readout Complete Audio
Notifier Widget starts automatically when you
select the Start Up ALL LRIS Control item
from the pulldown menu, and allows you to select the
sound you'd like to hear when readout has completed.
LRIS Slitmask Holder To Be Re-Designed (7 Oct 1998)
A design effort is underway to replace the existing
LRIS slitmask crossbar constraint with a holder that
mounts to the side of the slitmask frame, thereby
eliminating the crossbar as an obstruction to object
slits. An accessory to this holder will be a new
pickoff mirror mount that will bolt to the side of
the slitmask frame, extending into the beam with an
arm the approximate cross section of the existing
crossbar. This replacement hardware provides an
opportunity to increase the available area for the
pickoff mirror field, up to a factor of roughly 2,
by extending it in the LRIS field Y-axis into the
field visible in the guider on the other side of the
existing bar. The new pickoff mirror mount will be
fixed and should eliminate the current problems of
rotation and vignetting of the beam.
So, this is a poll to find out if users would benefit
from an increased pickoff field size, and whether we
need be very particular about the size and dimensions of
the field; or, are content with the existing size (35
arcsec square) because more area is available for slit
placement (indeed object field area increases with this
mount even with the same reflecting field, because the
existing pickoff mirror mount and crossbar will
disappear). Also, if we do keep the existing field
size, would it be acceptable to offset it to center in
the slitmask frame? In any case note that the slitmask
design programs will need to be modified to map the new
mirror and mount obstruction, and eliminate the
Thanks in advance for your input, and feel free to
forward this note to any LRIS multi-object observers
that I may have missed in the list.
LRIS Shutter Problems (15 Sep 1998)
The LRIS shutter control system has been returning
error conditions intermittently, causing CCD exposures
to fail. Inspection of the shutter mechanism suggests
that it is working fine, and that the problem is a
faulty limit switch. We have temporarily disabled the
circuit which returns the error code to the LRIS
control software. Although we expect no problems,
observers should be on the lookout for any symptoms of
shutter malfunction. Failure to open is indicated by
all exposures being ``dark,'' while failure to close
would manifest itself as a strong gradient in the
background level from the top of the chip to the
bottom, as well as pronounced vertical streaking in
the resultant image. The shutter limit switches will
be checked when LRIS next comes off the telescope in
Update: The shutter mechanism failed due to a broken
part on 17 Sep, and the broken part was fixed, restoring
``copymon'' Script Provides Automated Image Copies
(5 Aug 1998)
The new script copymon available on pelekane
will automatically copy incoming LRIS images from the
datataking system's output directory into another
directory for quick reduction. The user can start
this script running in the afternoon and allow it to
run all night (it terminates automatically after 10
a.m.). This frees the user to worry about reducing
the images rather than having to manually check for
new incoming images and copy them over. Please see
the help page for copymon
or consult your Instrument Specialist for further
Observer Accounts Enhanced
(31 Jul 1998)
In response to observer requests, the following enhancements
have been made to the numbered LRIS accounts in Waimea:
- The default login shell was changed from csh to
- The default login windows have been changed from
shelltool and cmdtool to
xgterm type, allowing full support for
tcsh command-line editing.
- IRAF and IDL are now accessible from the top-level
menu on the window manager. For reasons of setting
environment variables properly, use of the pulldown menu
to launch IDL and IRAF is preferred over typing
cl or idl on the command line.
- IRAF has been upgraded to allow use of the
cl or ncl command shells, and the
Ximtool, SAOtng, or
SAOimage display tools on all HQ Unix
machines (SunOS and Solaris).
- Startup of IRAF has been simplified by a script which
launches both the command shell of choice and the image
display tool of choice.
- The TAPE environment variable is set to the Exabyte
drive name if one exists on the machine to which you log
in, enabling easier operation of the tape drives.
- The most recent TeX/LaTeX/BibTeX and friends
(including xdvi) are now available on Solaris machines
so you can write up and submit your results before going
to bed in the morning.
LRIS Grating Turret Troubleshooting Guide
(23 Jul 1998)
- John Cromer has now provided three new macros to
help LRIS users diagnose and fix the most common causes of
LRIS grating move failures. These are documented in the
LRIS Grating Turret Troubleshooting
Guide. Please consult your Instrument Specialist
before attempting to execute these macros.
Motor Troubleshooting Guide
(05 Jul 1998)
- John Cromer's LRIS Motor Control Software Troubleshooting Guide is
now available on-line. This is really meant for use
by the Instrument Specialists and other ``black-belt'' LRIS
users --- please don't use this unless you really know what
you're doing. Some of the low-level commands lack
safeguards and can potentially damage the spectrograph.
Improve Slitmask Alignment Efficiency with Zeroth
Order Imaging (09 June 1998)
- Observers have reported success performing slitmask
aligning using zeroth-order imaging with the gratings. This
can save nearly 8 minutes per mask over the traditional method
of inserting the mirror to take mask alignment images. See
for more information.